OpenStudy (anonymous):

2x^4-162 divided by 2 PLEASE SOMEONE HELP

OpenStudy (anonymous):

Does it look like this? \[\frac{ 2x^4-162 }{ 2 }\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Okay, so I'm assuming we're trying to simplify this? So, 2/2= 1, so this will make it 1x^4 and -162/2 = -81 Can you figure out what your answer is?

OpenStudy (anonymous):

well i got on my test 2(x^4-81) and the teacher said to simplfy it more

OpenStudy (anonymous):

then i did 2(x^2-9)(x^2+9) and i dont know what to do afterwards it says to simplyfy it more

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Okay, so we're going to distribute. 2 * x^2, 2*-9, etc. If you do this and add the like terms together you should get the simplified version.

OpenStudy (whpalmer4):

The original is \[x^4-81\] after dividing through by 2 which seems plenty simplified to me. Maybe you were supposed to factor it? It's of the form \[a^2-b^2\] which factors to \[(a+b)(a-b)\] so we can write \[(x^2-9)(x^2+9)\] and we can factor both of those giving \[(x+3)(x-3)(x+3i)(x-3i)\]